OpenStudy (anonymous):

Find the gross income for each problem. Show all of your work. 44 hours at $10 per hour, with time and a half for over 40 hours

OpenStudy (anonymous):

40*10=400 4(10+5)=60 400+60=$460

OpenStudy (anonymous):

Thank you thank you! :)

OpenStudy (anonymous):

np